OpenStudy (anonymous):

If cotx=5/7,the ^2 is equivalent to what? How do i find quivalent to? a.)12/7/5 c)5sqrt(74 d)sqrt74

OpenStudy (anonymous):

square both sides. if cot(x)=5/7, squaring both sides gives [cot(x)]^2=(5/7)^2.

OpenStudy (anonymous):

or do you mean the square root?

OpenStudy (anonymous):

the sqrt

OpenStudy (anonymous):

makes more sense. do the same operation as I wrote earlier, just \[\sqrt{\cot(x)}=\sqrt{\frac{ 5 }{ 7}}\]

OpenStudy (anonymous):

and rationalize.